# An operational approach to classifying measurement incompatibility

### v1

@@ Abstract
Measurement incompatibility has proved to be an important resource for information processing tasks. In this work, we analyze various levels of incompatibility of measurement sets. We provide operational classification of measurement incompatibility with respect to two elementary classical operations, viz., coarse-graining of measurement outcomes and convex mixing of different measurements. We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or convex mixing. Robustness against white noise is investigated for mutually unbiased bases that can sustain full incompatibility. Furthermore, we propose operational witnesses for different levels of incompatibility subject to classical operations, using the input-output statistics of Bell-type experiments as well as experiments in the prepare-and-measure scenario.
@@
Measurement incompatibility is a concept relating observables which cannot be measured jointly with arbitrary accuracy Gühne et al. (2023). It is a purely quantum effect of which the most well-known example concerns the position and momentum of a quantum particle. Being a fundamental concept of quantum theory, it takes a pivotal role in explaining several quantum phenomena, such as, Bell-nonlocality Fine (1982a); Wolf et al. (2009), Einstein–Podolsky–Rosen steering Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015), measurement uncertainty relations Pramanik et al. (2013); Busch et al. (2014); Maity et al. (2017); Saha et al. (2020), quantum contextuality Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), quantum violation of macrorealism Mal and Majumdar (2016); Uola et al. (2019), and temporal and channel steering Karthik et al. (2015); Banik et al. (2015); Uola et al. (2018). Measurement incompatibility finds applications in various quantum information processing tasks. Recently, it has been shown that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al. (2023).
The significance of measurement incompatibility in various quantum effects calls for its in-depth characterization. Towards this direction, a classification of measurement incompatibility with respect to projection onto subspaces has been recently performed Uola et al. (2021). In the present work, our aim is to classify measurement incompatibility with respect to certain basic classical operations. From an operational perspective, here we employ two such operations, namely coarse-graining of measurement outcomes and convex mixing of measurement settings. Coarse-graining of measurement outcomes arises naturally Toscano et al. (2018) in several instances, such as, for example, quite obviously in measurements on continuous variable system. Though the eigen spectra of the observables are infinite-dimensional taking continuous values, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values. This inaccuracy in the recording of measurement outcomes is manifested in the coarse-graining of measurement outcomes, which is inevitable in practice. Coarse-graining has also been employed to study the phenomenon of quantum to classical transition. It is observed that quantum phenomena may disappear due to imprecision of measurement outcomes Kofler and Brukner (2007); Rudnicki et al. (2012); Jeong et al. (2014); Mal and Majumdar (2016).
@@
Device imperfection may also lead to the measurement device performing probabilistically a set of measurements, instead of always performing the desired particular measurement. In such a case, a convex mixing of the given set of measurements arises effectively. As incompatibility of measurements is a quantum concept, it is interesting to examine how this property behaves under such elementary classical operations. Our study is motivated by the question as to how one may compare the degree of incompatibility between two different sets of measurements. For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one. A similar argument holds for the case of convex-mixing of measurements too.
As incompatible measurements are useful for performing various information processing tasks, given a measurement device claiming to produce incompatible measurements, we must certify it before using it in an information processing task. Verification of incompatiblity of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning. Device-independent protocols are conceptually most powerful, relying only on the input-output statistics Brunner et al. (2014), with a wide range of applications from quantum cryptography Ekert (1991); Acín et al. (2007) to communication complexity Buhrman et al. (2010), self-testing Mayers and Yao (2004); Šupić and Bowles (2020); Goswami et al. (2018); Bian et al. (2020) and incompatibility witnesses Quintino et al. (2019). However, they require shared entanglement, an expensive resource, and prohibition of communication between the involved parties. On the other hand, semi-device-independent protocols are inspired by the standard prepare-and-measure scenario, with an additional constraint on the dimension of the communicated quantum states Van Himbeeck et al. (2017); Moreno et al. (2021). They have been applied in quantum key distribution Pawłowski and Brunner (2011), quantum networks Bowles et al. (2015), randomness certification Passaro et al. (2015), random access codes Li et al. (2012), and certification of meassurements Maity et al. (2021); Das et al. (2022).
In the present work we investigate the issue of certification of measurement incompatibility subjected to the classical operations of coarse-graining and convex mixing of measurements from both the device-independent and the semi-device-independent perspective. An essential ingredient in the performance of quantum information protocols is the presence of noise. Within the context of our present study, noise is reflected in the degrading of incompatibility properties of various measurement sets. We examine the critical amount of robustness of the noisy measurement above which the measurements remain incompatible under the mentioned classical operations. We then study incompatibility witnesses in the device-independent and the semi-device-independent scenarios.
@@
Our paper is structured as follows. In Sec. (II) we present the basic mathematical ingredients for classifying measurement incompatibility under the two operations of coarse-graining of outcomes and convex-mixing of measurements, respectively. In Sec. (III) we study the robustness under noise for various levels of incompatibility subjected to the above classical operations. In Sec. (IV) we define operational witnesses of incompatibility, and furnish examples to study their performance in both device-independent and the semi-device-independent frameworks. Concluding remarks are presented in Sec. (V).
II Classifying measurement incompatibility under elementary classical operations
@@
In the case of projective measurements, the observables whose corresponding operators commute are jointly measurable. However, this is not true in the case of general quantum measurements. The most general quantum measurement is described by Positive Operator Valued Measures (POVM), which is a set of operators, $\{O_{\alpha}\}$ with $0\leq O_{\alpha}\leq\mathbbm{1}$, and $\sum_{\alpha}O_{\alpha}=\mathbbm{1}$. A set of measurements defined by $\{M_{z_{x}|x}\}$ (here $x$ corresponds to the measurement, $z_{x}$ denotes the outcomes of the measurement labeled by $x$) is compatible if there exists a parent POVM, $G_{\lambda}$ and classical post-processing for each $x$ given by $\{p(z_{x}|x,\lambda)\}$ such that
$\forall z_{x},x,\quad M_{z_{x}|x}=\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda},$ (1)
where $0\leq p(z_{x}|x,\lambda)\leq 1,\sum_{z_{x}}p(z_{x}|x,\lambda)=1\forall x,\lambda$    Gühne et al. (2023).
II.1 Coarse-graining of outcomes
Coarse-graining of the outcomes of a measurement is a concept where some of the outcomes are treated on the same footing, i.e., different outcomes are clubbed together to form a single outcome. After the coarse-graining of the outcomes, the effective number of outcomes reduces. Coarse-graining may occur in practice due to measurement errors. Additionally, coarse-graining is used in when one does not require a finer description of the values of the measurement outcomes, but rather a broad description suffices to understand the nature of some physical property.
Let us introduce the notation $[k]:=\{0,\cdots,k-1\}$ for any natural number $k.$ Consider a $\overline{d}$ outcome measurement $\{M_{\overline{z}}\}$ and an arbitrary coarse-graining that yields a $d$ outcome measurement $\{M_{z}\}$, where $\overline{z}\in[\overline{d}]\,\,\text{and}\,\,z\in[d]$. The coarse-graining can be defined by conditional probabilities $\{c(z|\overline{z})\}$ such that
$M_{z}=\sum_{\overline{z}}c(z|\overline{z})M_{\overline{z}},$ (2)
and $c(z|\overline{z})\in\{0,1\}$. Trivial coarse-graining is the one where the measurement becomes single outcome, that is, $c(z|\overline{z})=1,\forall\overline{z}$. Two incompatible measurements may not necessarily remain incompatible after certain coarse-graining of their outcomes.
Definition 1 (Fully incompatible measurements w.r.t coarse-graining).

### v2

@@ Abstract
Measurement incompatibility has proved to be an important resource for information processing. In this work, we present an operational approach that leverages classical operations on the inputs and outputs of measurement devices to explore different layers of incompatibility among the measurements performed by the device. We study classifications of measurement incompatibility with respect to two types of classical operations, viz., coarse-graining of measurement outcomes and disjoint-convex-mixing of different measurements. We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or disjoint-convex-mixing. Robustness against white noise for different layers of incompatibility for mutually unbiased bases is investigated. Furthermore, we study operational witnesses for incompatibility subject to these classical operations, using the input-output statistics of Bell-type experiments as well as for experiments in the prepare-and-measure scenario.
@@
Measurement incompatibility is a concept related to observables that cannot be measured jointly with arbitrary accuracy Gühne et al. (2023). It is purely a quantum effect, of which the most well-known example concerns the position and momentum of a quantum particle. Being a fundamental concept of quantum theory, it takes a pivotal role in explaining several quantum phenomena, such as, Bell-nonlocality Fine (1982a); Wolf et al. (2009a), Einstein-Podolsky-Rosen steering Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015), measurement uncertainty relations Pramanik et al. (2013); Busch et al. (2014); Maity et al. (2017); Saha et al. (2020), quantum contextuality Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), quantum violation of macrorealism Mal and Majumdar (2016); Uola et al. (2019), and temporal and channel steering Karthik et al. (2015); Banik et al. (2015); Uola et al. (2018). Apart from the foundational significance of incompatible measurements, they have been proven to be an important resource for various information processing tasks Skrzypczyk et al. (2019); Buscemi et al. (2020); Carmeli et al. (2019). Recently, It is also known that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al. (2023).
The significance of measurement incompatibility in various operational tasks calls for its in-depth characterization. Towards this direction, a classification of measurement incompatibility with respect to projection onto subspaces has been recently performed Uola et al. (2021), also different hierarchy of incompatible measurements has been given based on the number of copies of quantum states required Carmeli et al. (2016), and the simulability of their statistics produced in the experiments Filippov et al. (2018); Egelhaaf et al. (2024). In the continuous variable system, the verification of incompatibility using phase-space quasiprobability distributions has been studied in Ref.Rahimi-Keshari et al. (2021).
In the present work, our objective is to classify measurement incompatibility in an operational approach that does not involve the details of a theory. To put it differently, we address how various degrees of incompatibility can be assessed solely by executing basic classical operations on the inputs or outputs of these measurements. From an operational perspective, when the concerned measurement devices are black boxes with no control over the internal workings, one can still realize different measurements by performing suitable classical operations to manipulate the statistics of the measurements. Here, we employ two such operations: coarse-graining of measurement outcomes (classical post-processing performed on the outcomes of a measurement) and convex-mixing of measurement settings (classical operation performed on the inputs). One can also consider classical operations performed on both inputs and outputs. Similar types of classical operations were also introduced in Haapasalo et al. (2012), nevertheless, the study of incompatibility under these classical operations was lacking. The motivation of this present work is to fill this crucial gap in the literature.
Coarse-graining of measurement outcomes arises naturally in several instances, for example, in measurements on continuous variable systems Toscano et al. (2018). Though the eigen spectra of the observables are infinite-dimensional and continuous, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values. This inaccuracy in the recording of measurement outcomes is manifested in the coarse-graining of measurement outcomes, which is inevitable in practice. On the other hand, device imperfection may also lead to the measurement device performing a set of measurements probabilistically, instead of always performing the desired particular measurement. In such a case, a convex-mixing of the given set of measurements arises effectively D’Ariano et al. (2005).
@@
By leveraging these classical operations on the measurement device, we can learn the finer details about the measurements i.e. the different layers of incompatibility, which can give a measure of the degree of incompatibility of a set of measurements. This can set a benchmark for legitimately choosing incompatible measurements for an information processing task. Our study is motivated to address how one may compare the degree of incompatibility between two different sets of measurements subjected to the aforementioned classical operations. For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one. A similar argument holds for the case of convex-mixing of measurements also.
In this work, we establish analytical criteria for determining when a pair of projective measurements are fully incompatible, i.e., remain incompatible under all possible coarse-grainings of measurement outcomes. We further analyze the full incompatibility of a set of three qubit measurements under all possible convex-mixing as well. Within the context of our present study, noise is reflected in degrading the incompatibility properties of various measurement sets Designolle et al. (2019a). We compute the critical noise threshold below which the mutually unbiased bases measurements remain incompatible under the above-mentioned classical operations.
As incompatible measurements are useful for various information processing tasks, any device claiming to produce incompatible measurements must be certified before using it in an experiment. Verification of incompatibility of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning. This can be done in two ways one is a device-independent way utilizing Bell-type experiments Wolf et al. (2009b); Bene and Vértesi (2018); Hirsch et al. (2018) and another is a semi-device-independent way inspired by the standard prepare-and-measure scenario de Gois et al. (2021); Saha et al. (2023); Carmeli et al. (2020). In the present work, we investigate the issue of certification of different layers of measurement incompatibility under the introduced classical operations both from the device-independent and the semi-device-independent perspective. As both Bell-type experiments and prepare-and-measure experiments Zhang et al. (2024); Tavakoli et al. (2015) are viable with the current technology, by performing our introduced classical operations on those experiments we can operationally certify the different layers of incompatibility thus allowing us to know the subtleties of incompatibility about the measurement device which is a new and interesting offshoot of our work.
@@
Our paper is structured as follows. In Sec.(II), the significance of these classical operations in an operational paradigm is discussed along with the physical ground for these classical operations. In Sec. (III), we define a hierarchy in the incompatibility of measurements under these classical operations. In Sec. (IV), we study the noise robustness for various levels of incompatibility of measurements subjected to the above classical operations. In Sec. (V), we define operational witnesses of incompatibility of measurements under these classical operations and furnish examples to study their performance in device-independent and semi-device-independent frameworks. Concluding remarks are presented in Sec.(VI).
II Operational approach to classifying incompatibility
Consider a measurement device where, when a physical system is probed, we can choose measurement settings via a classical parameter, denoted by $x$. This parameter serves as the input to the device, and the device produces the measurement outcome $z$, corresponding to that input. Suppose there are $n$ possible inputs, i.e., $x\in[n]$, and for each input, there are $d$ possible outcomes, i.e., $z\in[d]$, where $[k]$ represents the set $\{0,\cdots,k-1\}$ for any natural number $k$. For simplicity, we assume that the number of outcomes for each measurement is the same. Without loss of generality, this assumption holds, as any measurement with fewer than $d$ outcomes can be treated as having $d$ outcomes by considering some outcomes as never occurring. By probing different system preparations, we can gather input-output statistics from the device.
@@ II Operational approach to classifying incompatibility
In quantum mechanics, measurements are described by Positive Operator Valued Measures (POVMs), which consist of positive semi-definite operators that sum up to the identity operator. Let us represent the measurements realized in the device by the set of operators $\{M_{z|x}\}_{z,x}$, where $x$ indexes different measurements, and $z$ denotes the corresponding outcomes. A set of measurements is said to be compatible if there exists a parent POVM, $G_{\lambda}$, and classical post-processing $\{p(z|x,\lambda)\}$ for each $x$ such that
$\forall z,x,\quad M_{z|x}=\sum_{\lambda}p(z|x,\lambda)G_{\lambda},$ (1)
where $0\leq p(z|x,\lambda)\leq 1,$ and $\sum_{z}p(z|x,\lambda)=1,$ for all $x,\lambda$ Gühne et al. (2023). As a special case, if the operators are projectors, then the two measurements are jointly measurable when their corresponding operators commute.
In an operational paradigm, we do not have direct control over the internal workings of the measurement device. However, we can manipulate the classical inputs and outputs to realize different measurements.
Figure 1: Classical operations on the outputs (left) and classical operations on the inputs (right) to realize a new set of measurements are depicted.
II.1 Classical operations on outputs
In general, one can modify the outcomes of a measurement by applying a post-processing operation or stochastic map (see FIG. 1). For a given input $x$, this operation is mathematically described by a probability distribution $\mathcal{C}_{x}:=\{c_{x}(\tilde{z}_{x}|z)\}$, $c_{x}(\tilde{z}_{x}|z)\geqslant 0$ and $\sum_{\tilde{z}_{x}}c_{x}(\tilde{z}_{x}|z)=1$. Here $\tilde{z}_{x}\in[\tilde{d}_{x}]$ denotes the outcome of the new measurement after the operation, and $\mathcal{C}_{x}$ is the stochastic map with $c_{x}(\tilde{z}_{x}|z)$ is the probability of producing outcome $\tilde{z}_{x}$ when outcome $z$ is obtained for the measurement $x$. In general, one can apply different operations for different $x$. By applying different operations for each $x$, one can generate a new set of measurements described by the operators $\{\tilde{M}_{\tilde{z}_{x}|x}\}_{\tilde{z}_{x},x}$ as follows:
$\tilde{M}_{\tilde{z}_{x}|x}=\sum_{z}c_{x}(\tilde{z}_{x}|z)M_{z|x}\,\,,$ (2)
for each $x.$ If the initial set of measurements is incompatible, we can ask whether this new set is incompatible or not.
The set of all possible operations $\{\mathcal{C}_{x}\}$ is infinite. Furthermore, certain operations, such as generating random outputs, will always result in a set of measurements that is compatible, regardless of the original measurements. To avoid such intricacies, we focus on a particular class of operations where $c_{x}(\tilde{z}_{x}|z)\in\{0,1\}$ and $\tilde{d}_{x}<d$. These classical operations involve coarse-graining or relabelling/permuting outcomes (or a combination of both). This class contains a finite number of operations, and interestingly, any classical operation on outputs can be expressed as convex combinations of this class of operations. More importantly, certain quantum measurements remain incompatible even after all possible operations of this kind. We will refer to this class of operations as coarse-graining. Coarse-graining typically refers to treating more than one outcome as equivalent, i.e., multiple outcomes are clubbed together to form a single outcome. As a result, the effective number of outcomes is reduced.
II.2 Classical operation on inputs
We can also apply classical operations to the inputs to generate a new set of measurements, each of which is a convex combination of the initial measurements (see FIG. 1). This classical operation is represented by a probability distribution $\{e(x|\tilde{x})\}_{x,\tilde{x}}$, where $e(x|\tilde{x})\geqslant 0$ and $\sum_{x}e(x|\tilde{x})=1$ for each $\tilde{x}\in[m]$. Here, $e(x|\tilde{x})$ refers to the convex weightage of appearing measurement $x$ in the new measurement labelled by $\tilde{x}$. The new set of $m$ measurements, defined by the operators $\{\tilde{M}_{z|\tilde{x}}\}_{z,\tilde{x}}$, is given by:
$\tilde{M}_{z|\tilde{x}}=\sum_{x}e(x|\tilde{x})M_{z|x},$ (3)
for every $z$.
If the same convex mixing is applied uniformly to all inputs, i.e., $e(x|\tilde{x})$ is the same for all $\tilde{x}$, the resulting measurements will be identical and hence compatible. To avoid such cases, we introduce a particular class of convex mixing, which we will refer to as disjoint-convex-mixing of measurements. In disjoint-convex-mixing, no initial measurement appears in the convex combination of more than one new measurement. Mathematically, for every $z$, only one of the numbers from the set $\{e(x|\tilde{x})\}_{\tilde{x}}$ is non-zero, and others are zero, which can be expressed as:
$\forall x,\ \max_{\tilde{x}}\{e(x|\tilde{x})\}=\sum_{\tilde{x}}e(x|\tilde{x}).$ (4)
In this work, we focus on disjoint-convex-mixing for classifying measurement incompatibility. We will point out later that there are measurements that remain incompatible after all possible disjoint-convex-mixing.
II.3 Classical operations on both inputs and outputs
It is also possible to construct new measurements by combining operations on both inputs and outputs. This can be done in two ways - first applying an operation on the inputs followed by the outputs, or vice versa.
In the former case, the operators $\{\tilde{M}_{\tilde{z}_{\tilde{x}}|\tilde{x}}\}_{\tilde{z}_{\tilde{x}},\tilde{x}}$ defining the new measurements are
$\tilde{M}_{\tilde{z}_{\tilde{x}}|\tilde{x}}=\sum_{z}\sum_{x}c(\tilde{z}_{x}|z)e(x|\tilde{x})M_{z|x}\ ,$ (5)
where $\{e(x|\tilde{x})\}$ and $\{c(\tilde{z}_{x}|z)\}_{x}$ denote the classical operation on inputs and outputs, respectively. In the latter case, the set of operators $\{\tilde{M}_{\tilde{z}|\tilde{x}}\}$ are
$\tilde{M}_{\tilde{z}|\tilde{x}}=\sum_{x}\sum_{z}e(x|\tilde{x})c(\tilde{z}|z)M_{z|x}\ ,$ (6)
where $\{e(x|\tilde{x})\}$ and $\{c(\tilde{z}|z)\}$ denote the classical operation on inputs and outputs, respectively. Here again, without loss of generality, we consider the number of outcomes after the operations on the outcomes to be the same, which is denoted by $\tilde{z}$.
It is important to note that these two cases are not equivalent, as the stochastic operations on the outputs apply to different sets of measurements in each case.
In summary, one can explore the effects of various classical operations on both inputs and outputs to determine whether the resulting measurements remain incompatible. However, this work specifically focuses on two key classes of operations: coarse-graining of outputs and disjoint-convex-mixing of inputs. In addition to the operational relevance of the two types of operations mentioned above, their significance for practical implementation is discussed in the following subsection.
II.4 Physical motivations for considering the two specific classical operations
Let us now elaborate further the motivations for considering the two aforementioned specific classical operations on a set of measurements. Coarse-graining of measurement outcomes is a natural consequence of an observer’s limitation in a practical scenario involving multi-outcome measurements. For example, consider the measurement of the spin-$z$ component (associated with the operator $J_{z}$) of the spin-$j$ system, where $j$ is very large. The possible outcomes of this measurement are nothing but the eigenvalues of $J_{z}$, denoted by $m\in\{-j,-j+1,\cdots,j\}$. In practice, this type of measurement is performed using a Stern-Gerlach type experiment. In such a practical situation, the concept of “neighbouring outcomes” arises Kofler and Brukner (2007). For example, the outcomes $m$ and $m+1$ are neighbouring outcomes in the real configuration space in a Stern-Gerlach type experiment. For large $j$, it is almost impossible for a device with finite precision to resolve these neighbouring outcomes in the observation screen – giving rise to coarse-graining of measurement outcomes. This type of practical limitation is not only limited to large spins but also applicable to other multi-outcome measurements and measurements of continuous variables (e.g., position, momentum Das et al. (2024)). Note that coarse-graining of measurement outcomes has been invoked in the context of explaining classical limits of quantum mechanics Kofler and Brukner (2007); Rudnicki et al. (2012); Jeong et al. (2014); Mal and Majumdar (2016); Mal et al. (2016); Mukherjee et al. (2019).
On the other hand, convex mixing of measurements is also inevitable in a practical situation. To explain it in more detail, let us consider the example of measuring $\sigma_{z}$ (Pauli spin observable) on a qubit. In practice, the observable to be measured ($\sigma_{z}$ in the present example), defined by the relative direction of the inhomogeneous magnetic field in the Stern-Gerlach apparatus with respect to the direction of the incoming beam of spin-$1/2$ particles, may not be kept fixed in all experimental runs. Consequently, instead of $\sigma_{z}$, $\vec{\sigma}.\hat{n}$ will be measured ($\vec{\sigma}.\hat{n}$ = $\sigma_{x}n_{x}+\sigma_{y}n_{y}+\sigma_{z}n_{z}$ with $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ are three Pauli operators and $n_{x}^{2}+n_{y}^{2}+n_{z}^{2}=1$), where $\hat{n}$ will be different in different runs (all such $\hat{n}$ should be close to $\hat{z}$). Effectively, this will give rise to convex mixing of different measurements.
Hence, these operations are extremely relevant in practical scenarios. The various layers of incompatibility and their operational implications under these operations are discussed in the subsequent sections.
III Classifying measurement incompatibility under coarse-graining and disjoint-convex-mixing
III.1 Coarse-graining of outcomes
Recall that the coarse-graining of a set of $d$-outcome measurements $\{M_{z|x}\}$ produces new measurements defined by
$\tilde{M}_{\tilde{z}_{x}|x}=\sum_{z}c_{x}(\tilde{z}_{x}|z)M_{z|x},\ \text{ where }\ c_{x}(\tilde{z}_{x}|z)\in\{0,1\}.$ (7)
Before moving forward, let us define what we mean by trivial coarse-graining. If the coarse-graining results in a measurement where one of the outcomes, say $\tilde{z}_{x}$, always occurs, that is, $\tilde{M}_{\tilde{z}_{x}|x}=\mathbbm{1}$, we refer that operation as trivial coarse-graining for the measurement labelled by $x$. The fact that two incompatible measurements may not necessarily remain incompatible after certain coarse-graining motivates the following definition of full incompatibility with respect to (w.r.t.) coarse-graining.
Definition 1 (Fully incompatible measurements w.r.t. coarse-graining).